Calculate the probability of winning the indicated lottery.

Pennsylvania Lucky for life lotto: Select the six winning numbers from 1, 2, ........... 38.

Texas cash five: Select the five winning numbers from 1, 2, ........ 37
Answer #1

Pennsylvania Lucky for life lotto

Total different numbers : 38

Lottery winning combination : 1

Combinations of all 6 numbers : 38 C 6

Probability ( Win) = 1 / 38 C 6   = 1 / 2760681

Texas cash five

Total different numbers : 37

Lottery winning combination : 1

Combinations of all 5 numbers : 37C 5

Probability ( Win) = 1 / 37 C 5

  = 1/ 435897